Data from a small bookstore are shown in the accompanying table. The manager wants to predict Sales from Number of Sales People Working. Complete parts a through h below.

Number of sales people working

Sales​ (in $1000)

3

10

4

11

5

13

9

15

11

18

11

20

12

20

14

23

17

23

20

25

x overbar x=10.6

y overbar y=17.8

​SD(x)=5.56

​SD(y)=5.31

​a) Find the slope​ estimate, b1.

b1=???? ​(Round to two decimal places as​ needed.)
Answer #1

Solution:

The formula for slope is given as below:

Slope = b1 = r*(Sy/Sx)

We are given Sy = 5.31

And Sx = 5.56

Now, we have to find the value for the correlation coefficient r. The formula for the correlation coefficient is given as below:

Correlation coefficient = r = [n∑xy - ∑x∑y]/sqrt[(n∑x^2 – (∑x)^2)*(n∑y^2 – (∑y)^2)]

The calculation table for the correlation coefficient is given as below:

.S no. X Y XY X^2 Y^2
1 3 10 30 9 100
2 4 11 44 16 121
3 5 13 65 25 169
4 9 15 135 81 225
5 11 18 198 121 324
6 11 20 220 121 400
7 12 20 240 144 400
8 14 23 322 196 529
9 17 23 391 289 529
10 20 25 500 400 625
Total 106 178 2145 1402 3422


Correlation coefficient = r = 0.971733 Correlation coefficient = r = [10*2145 – 106*178]/sqrt[(10*1402– 106*106)*(10*3422 – 178*178)]

Slope = b1 = r*(Sy/Sx)

Slope = b1 = 0.971733*(5.31/5.56) = 0.928039

Required answer: b1 = 0.928039